EVALUATING RISK AND RETURN

Stock X has a 9.5% expected return, a beta coefficient of 0.8, and a 35% standard deviation of expected returns. Stock Y has a 12.5% expected return, a beta coefficient of 1.2, and a 20.0% standard deviation. The risk-free rate is 6%, and the market risk premium is 5%.

Calculate each stock's coefficient of variation. Round your answers to two decimal places. Do not round intermediate calculations.

CVx =

CVy =
Which stock is riskier for a diversified investor?

For diversified investors the relevant risk is measured by standard deviation of expected returns. Therefore, the stock with the higher standard deviation of expected returns is more risky. Stock X has the higher standard deviation so it is more risky than Stock Y.
For diversified investors the relevant risk is measured by beta. Therefore, the stock with the lower beta is more risky. Stock X has the lower beta so it is more risky than Stock Y.
For diversified investors the relevant risk is measured by standard deviation of expected returns. Therefore, the stock with the lower standard deviation of expected returns is more risky. Stock Y has the lower standard deviation so it is more risky than Stock X.
For diversified investors the relevant risk is measured by beta. Therefore, the stock with the higher beta is less risky. Stock Y has the higher beta so it is less risky than Stock X.
For diversified investors the relevant risk is measured by beta. Therefore, the stock with the higher beta is more risky. Stock Y has the higher beta so it is more risky than Stock X.

Calculate each stock's required rate of return. Round your answers to two decimal places.

rx = %

ry = %
On the basis of the two stocks' expected and required returns, which stock would be more attractive to a diversified investor?

Calculate the required return of a portfolio that has $8,500 invested in Stock X and $2,500 invested in Stock Y. Do not round intermediate calculations. Round your answer to two decimal places.

rp = %

If the market risk premium increased to 6%, which of the two stocks would have the larger increase in its required return?
Answer #1

CVx =Standard Deviation of X/Expected Return of Stock X =35%/9.5% =3.68
CVy =Standard Deviation of Y/Expected Return of Stock Y  =20%/12.5% =1.60

Option V For diversified investors the relevant risk is measured by beta. Therefore, the stock with the higher beta is more risky. Stock Y has the higher beta so it is more risky than Stock X. is correct option

Required Rate of X =Risk Free Rate+Beta*Market Risk Premium =6%+0.8*5% =10%
Required Rate of Y =Risk Free Rate+Beta*Market Risk Premium =6%+1.2*5% =12%

On the basis of expected rate and required rate Stock Y is more attractive as 12.5% is more than 12%.

Required Rate of Portfolio =Weight of X*Return of X+Weight of Y*Return of Y =8500/(8500+2500)*9.5%+2500/(8500+2500)*12.5% =10.18%

Since Stock Y has higher beta , Stock Y will have larger increase in required rate with increase in risk premium